OpenStudy (ny,ny):

M= Rise/ Run so if the slope is -2, the rise will be -2 and the run will be 1. Is the run positive or negative? I'm going to use this slope of an equation to graph and i want to know which direction the line will go.

OpenStudy (theeric):

Well, just algebraically you can see\[-2=-\frac{2}{1}=\frac{-2}{1}=\frac{2}{-1}\]

OpenStudy (theeric):

So if you have a negative rise (look at the third expression for an example) \[(\frac{-2}{1})\] then you have a positive 1 as your run.

OpenStudy (ny,ny):

oh, so if the rise is a whole number and it is negative, the run will be positive

OpenStudy (theeric):

As long as you let your rise be negative! :) You can actually work it out both ways, as long as just one of rise or run is negative. Look at -2/1. Then you go down 2, and right 1. Look at 2/-1. Then you go up 2, and left 1. It's just because they're opposite! -2 from 2, and 1 from -1, I mean.

OpenStudy (ny,ny):

yeah I get it :)

OpenStudy (ny,ny):

oh wait, say the slope is -2/3 then is the rise is negative and the run positive?

OpenStudy (theeric):